OpenStudy (anonymous):

sorry guys find the exact value of the expression log2 6-log2 15+log 20

OpenStudy (anonymous):

why you gave space between 2 and 6 does it mean it is multiplied with log2 plz use brackets

OpenStudy (campbell_st):

well\[\log _{2} 6 - \log _{2} 15 = \log _{2} (6/15)\] and \[\log _{2} 15 + \log _{2} 20 = \log _{2} (15\times20)\] hope this helps as I'm not sure what the actual question is

OpenStudy (campbell_st):

an alternative is \[\log _{2} (6/15\times20) = \log _{2} (8) = 3\]

OpenStudy (campbell_st):

the reason is 2^3 = 8

OpenStudy (anonymous):

plz dont give him one more answer @campbell_st

OpenStudy (campbell_st):

well I think the answer is 3...

OpenStudy (anonymous):

yes it is three thank you :)